b+b+b+7=b+22 I need the value of b

Answer by oberobic(2304)

b+b+b+7 = b+22
collect terms
3b + 7 = b + 22
subtract b from both sides
2b + 7 = 22
subtract 7 from both sides
2b = 15
divide both sides by 2
b = 15/2
b = 7 1/2